# Scale-Invariant Gradient Aggregation for Constrained Multi-Objective Reinforcement Learning

### v1

@@ 6 Experiments
This section evaluates the proposed method and baselines across various tasks with and without constraints. First, we explain how methods are evaluated on overall tasks and then present the CMORL baselines. Subsequently, each task is described, and the results are analyzed. Finally, we conclude this section with ablation studies of the proposed method.
Table 1: Results of the Safety Gymnasium tasks. The hypervolume (HV), sparsity (SP), and constraint values (Con1, Con2) are obtained by averaging results from five random seeds. The thresholds for each constraint are shown after Con1 and Con2, and elements that violate constraints are shown in red. Since the single-agent tasks have only a single constraint, so Con2 does not exist.
Single-Point-Goal Single-Car-Goal
HV (×$\times 10^{1}$) $\uparrow$ $\overline{\mathrm{SP}}$ (×$\times 10^{-1}$) $\downarrow$ Con1 ($\leq 0.025$) Con2 ($\leq 0.025$) HV (×$\times 10^{1}$) $\uparrow$ $\overline{\mathrm{SP}}$ (×$\times 10^{-1}$) $\downarrow$ Con1 ($\leq 0.025$) Con2 ($\leq 0.025$)
CAPQL $2.347$ $0.861$ $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}0.026$ $0.085$ $4.472$ $0.019$
PD-MORL $2.709$ $1.255$ $0.024$ $0.070$ $3.112$ $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}0.027$
LP3 $\mathbf{3.807}$ $4.441$ $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}0.035$ $0.701$ $15.519$ $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}0.037$
CoMOGA (Ours) $3.628$ $\mathbf{0.511}$ $0.022$ $\mathbf{1.057}$ $\mathbf{0.558}$ $0.024$
Multi-Point-Goal Multi-Car-Goal
CAPQL $17.089$ $\mathbf{1.445}$ $0.019$ $0.023$ $51.401$ $1.604$ $0.016$ $0.022$
@@
CoMOGA (Ours) $\mathbf{38.646}$ $2.010$ $0.022$ $0.019$ $\mathbf{65.895}$ $\mathbf{1.115}$ $0.022$ $0.017$
6.1 Evaluation Metrics and Baselines
Once the universal policy is trained, an estimated CP front, denoted as $P\subset\mathbb{R}^{N}$, can be obtained by calculating the objective values $(J_{R_{1}}(\pi),...,J_{R_{N}}(\pi))$ for a range of preferences. Given that CMORL aims to approximate the ground truth CP front, it is essential to assess how close the estimated CP front is to the ground truth, and it can be realized by measuring the coverage and density of the estimated front. To this end, we use the hypervolume (HV) for coverage measurement and sparsity (SP) for density estimation, which are commonly used in many MORL algorithms (Xu et al., 2020; Basaklar et al., 2023). Given an estimated CP front $P\subset\mathbb{R}^{N}$ and a reference point $r\in\mathbb{R}^{N}$, the hypervolume (HV) is defined as follows (Zitzler & Thiele, 1999):
[…]
where $Q(P,r):=\{z|\exists p\in P\;\text{s.t.}\;r\preceq z\preceq p\}$. HV is then the volume of the area surrounding the reference point and the estimated CP front. For a detailed explanation of this metric, including a visual description and how the reference point is set, please refer to Appendix B.2. Sparsity (SP), on the other hand, measures how uniformly the estimated CP fronts are distributed and is defined as the average of the squared distances between elements (Xu et al., 2020):
$\small\mathrm{SP}(P):=\frac{1}{|P|-1}\sum\nolimits_{j=1}^{N}\sum\nolimits_{i=1}^{|P|-1}(\tilde{P}_{j}(i)-\tilde{P}_{j}(i+1))^{2},$ (15)
where $\tilde{P}_{j}:=\mathrm{Sort}(\{p_{j}|\forall p\in P\})$, and $\tilde{P}_{j}(i)$ is the $i$th element in $\tilde{P}_{j}$. Since a larger value indicates a less dense set, it is preferred to have a small value. However, given that the cardinality of the set $P$ is constant, a larger HV typically corresponds to a greater SP. To remove this inherent correlation, we propose the following normalized version:
$\small\overline{\mathrm{SP}}(P):=\frac{1}{|P|-1}\sum_{j=1}^{N}\sum_{i=1}^{|P|-1}\left(\frac{\tilde{P}_{j}(i)-\tilde{P}_{j}(i+1)}{\mathrm{max}_{k}\tilde{P}_{j}(k)-\mathrm{min}_{k}\tilde{P}_{j}(k)}\right)^{2},$ (16)
where the elements of the sorted array $\tilde{P}_{j}$ are normalized with respect to their maximum and minimum values. By using the normalized sparsity, we can eliminate the correlation with HV while preserving the ability to measure sparsity.
For comparison with the proposed method, we first employ LP3 (Huang et al., 2022), a CMORL algorithm, as a baseline. This method extends the scale-invariant MORL algorithm, MOMPO (Abdolmaleki et al., 2020), to CMORL by handling constraints in a similar way to the Lagrangian method. Subsequently, we utilize state-of-the-art actor-critic-based MORL algorithms, PD-MORL (Basaklar et al., 2023) and CAPQL (Lu et al., 2023), which are based on linear scalarization. They are respectively implemented based on the TD3 (Fujimoto et al., 2018) and SAC (Haarnoja et al., 2018) algorithms. We extend them to CMORL algorithms by handling constraints using the Lagrangian method, like LP3, and please see Appendix B.3.1 for implementation details.
Table 2: Results of the locomotion tasks. The hypervolume (HV), sparsity (SP), and constraint values (Con1, Con2, Con3) are obtained by averaging results from five random seeds. The thresholds for each constraint are shown after Con1, Con2, and Con3, and elements that violate the constraint are shown in red.
Quadruped Biped
HV (×$\times 10^{4}$) $\uparrow$ $\overline{\mathrm{SP}}$ (×$\times 10^{-1}$) $\downarrow$ Con1 ($\leq 0.025$) Con2 ($\leq 0.025$) Con3 ($\leq 0.4$) HV (×$\times 10^{4}$) $\uparrow$ $\overline{\mathrm{SP}}$ (×$\times 10^{-1}$) $\downarrow$ Con1 ($\leq 0.025$) Con2 ($\leq 0.025$) Con3 ($\leq 0.4$)
[…]
LP3 $4.637$ $0.170$ $0.024$ $0.009$ $0.201$ $0.114$ $1.991$ $0.004$ $0.002$ $0.356$
CoMOGA (Ours) $\mathbf{4.936}$ $\mathbf{0.169}$ $0.015$ $0.014$ $0.338$ $\mathbf{7.067}$ $\mathbf{0.177}$ $0.002$ $0.002$ $0.359$
Table 3: Results of the MO-Gymnasium tasks. HV and SP are obtained by averaging results from five random seeds.
Half-Cheetah Hopper Water-Reservoir Lunar-Lander
HV (×$\times 10^{5}$) $\uparrow$ $\overline{\mathrm{SP}}$ (×$\times 10^{-1}$) $\downarrow$ HV (×$\times 10^{8}$) $\uparrow$ $\overline{\mathrm{SP}}$ (×$\times 10^{-1}$) $\downarrow$ HV (×$\times 10^{4}$) $\uparrow$ $\overline{\mathrm{SP}}$ (×$\times 10^{-1}$) $\downarrow$ HV (×$\times 10^{8}$) $\uparrow$ $\overline{\mathrm{SP}}$ (×$\times 10^{-1}$) $\downarrow$
CAPQL $10.418$ $1.211$ $0.152$ $8.103$ $1.869$ $0.688$ $13.517$ $\mathbf{0.180}$
PD-MORL $0.959$ $2.045$ $0.064$ $0.786$ $\mathbf{2.829}$ $1.022$ $9.379$ $3.419$
LP3 $3.557$ $0.720$ $0.000$ $0.877$ $1.417$ $1.615$ $0.606$ $9.471$
CoMOGA (Ours) $\mathbf{16.903}$ $\mathbf{0.166}$ $\mathbf{17.468}$ $\mathbf{0.492}$ $1.670$ $\mathbf{0.580}$ $\mathbf{17.365}$ $0.259$
6.2 Safety Gymnasium
We utilize single-agent and multi-agent goal tasks from the constrained RL environment, Safety Gymnasium (Ji et al., 2023). These tasks require navigating robots to designated goals while avoiding obstacles. The single-agent goal task has two objectives: 1) reaching goals as many times as possible within a time limit and 2) maximizing energy efficiency, along with a single constraint to avoid collisions with obstacles. Conversely, the multi-agent goal task involves two agents and two goals, requiring control of both agents to reach goals competitively. It encompasses two objectives, maximizing goal achievement for each agent, and two constraints to avoid obstacle collisions for each robot. Various types of robots are available for these tasks, among which point and car robots are used. For details on the tasks and hyperparameter settings, please refer to Appendix B.
The results are presented in Table 1, and the visualization of the estimated Pareto fronts can be found in Appendix C. The proposed method shows superior performance, achieving the highest HV and the lowest SP in three out of four tasks while satisfying the constraints across all tasks. On the other hand, other baselines violate constraints in at least one task, confirming that the proposed method can handle constraints more stably as it does not need additional optimization variables. LP3 shows the best HV in the Single-Point-Goal but violates the constraint. Moreover, in the Multi-Point-Goal, the SP of LP3 cannot be obtained because the estimated CP front of LP3 contains only one element due to constraint violation, so we leave this item empty in Table 1.
6.3 Legged Robot Locomotion
To evaluate the proposed method in more complex environments, we utilize the locomotion tasks from Kim et al. (2023), where the dimension of the action space is at least ten. The task is to control a quadrupedal or bipedal robot to follow randomly given commands while satisfying three constraints: keeping 1) body balance, 2) CoM height, and 3) pre-defined foot contact timing. Additionally, we have modified the tasks to have two objectives: 1) matching the current velocity with the command and 2) minimizing energy consumption. Please see Appendix B for more details.
We present the results of locomotion tasks in Table 2, and the proposed method shows the best HV and SP while satisfying constraints in all tasks. In these tasks, handling constraints using the Lagrangian method is more challenging since the number of Lagrange multipliers increases due to the multiple constraints. Because of this, PD-MORL and CAPQL show low HV and high SP values. We cannot calculate the SP of PD-MORL in the quadrupedal task for the same reason as in the Multi-Point-Goal task of LP3, so we leave it empty. LP3 shows comparable performance to the proposed method in the quadrupedal task, suggesting that LP3 can update policies more stably than other baselines because it updates the policy in the distribution space rather than the reward space.
6.4 Multi-Objective Gymnasium
We conduct experiments in the Multi-Objective (MO) Gymnasium (Alegre et al., 2022), which is a well-known MORL environment, to examine whether the proposed method also performs well on unconstrained MORL tasks. We use tasks with continuous action spaces in the MO Gymnasium, and there are four tasks available: Half-Cheetah, Hopper, Water-Reservoir, and Lunar-Lander. Each task has two, three, two, and four objectives, respectively. For more details, please refer to Appendix B. The results are presented in Table 3, and the proposed method achieves the best HV and SP in three out of the four tasks. These results suggest that the proposed gradient aggregation strategy, which involves transforming objectives into constraints, is also effective in unconstrained multi-objective tasks.
(a) Visualization of estimated Pareto fronts
(b) Performance metrics
Figure 3: Comparison CoMOGA with linear scalarization. The estimated CP fronts obtained from CoMOGA and LS are visualized in (a). For each scaled task, we conduct experiments with five random seeds, plotting the fronts from each seed in the same figure. (b) presents bar graphs showing HV and SP values for each scaled task, with scale details indicated on the x-axis.
6.5 Invariance to Objective Scales
In this section, we compare the proposed method with the linear scalarization approach in terms of sensitivity to the scale of objectives. To do this, we prepare a linear scalarization (LS) version of CoMOGA, which linearly sums the objectives instead of converting them into constraints. We modify the scale of the energy efficiency objective in the quadrupedal task by providing agents with 10×10$\times$10 × and 0.1×0.1$\times$0.1 × energy efficiency rewards. The results of the trained universal policy are visualized in Figure 3, and the values of the HV and SP metrics are reported in Appendix C. Observing Figure 3(a), CoMOGA exhibits consistent results between the original and scaled tasks. However, the results of the LS version are concentrated in the upper left for the 10×10$\times$10 × scaled task and in the lower right for the 0.1×0.1$\times$0.1 × scaled task, confirming that the LS approach tends to prioritize objectives with large scales. Additionally, Figure 3(b) shows that the metrics of CoMOGA do not vary with the objective scales, but the metrics of LS change significantly. In particular, the SP becomes more than three times larger in the scaled tasks compared to the original task. Through this study, we highlight the sensitivity of LS approaches to the objective scales and indicate that this sensitivity can lead to a degradation in performance.
7 Conclusions
We have introduced a CMORL algorithm, called CoMOGA, to maximize multiple objectives while satisfying constraints. The proposed method is based on a novel transformation process that converts objectives into constraints, which makes the method invariant to the objective scales and capable of handling constraints without additional optimization variables. We theoretically demonstrate that the proposed method converges to a local Pareto optimal policy and show that the proposed method outperforms the baseline methods in terms of CP front coverage and constraint satisfaction through various experiments.
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Theoretical Analysis
A.1 Transformation Process
In this section, we prove that the solutions of (5) and (6) are equivalent. Under the linear assumption, we can express the $i$th objective function as $g_{i}^{T}g+J_{R_{i}}(\theta_{\mathrm{old}})$. Then, (5) can be rewritten as follows:
$\mathrm{max}_{g}\;g_{i}^{T}g+J_{R_{i}}(\theta_{\mathrm{old}})\;\;\mathrm{s.t.}\;\sqrt{g^{T}Hg}\leq\epsilon.$ (17)
$\mathrm{max}_{\lambda\geq 0}\mathrm{min}_{g}\;-g_{i}^{T}g+\lambda(\sqrt{g^{T}Hg}-\epsilon)=:L(g,\lambda).$ (18)
Then, $g^{*}(\lambda)=\mathrm{argmin}_{g}L(g,\lambda)=\frac{1}{2\lambda}H^{-1}g_{i}$. By substituting $g^{*}(\lambda)$ to $g$, we get $\lambda^{*}=\mathrm{argmax}_{\lambda}L(g^{*},\lambda)=\frac{1}{2\epsilon}\sqrt{g_{i}^{T}H^{-1}g_{i}}$. Finally, we obtain the solution, $g^{*}=g^{*}(\lambda^{*})=\epsilon H^{-1}g_{i}/\sqrt{g_{i}^{T}H^{-1}g_{i}}$.
$\mathrm{min}_{g}\;g^{T}Hg\;\;\mathrm{s.t.}\;e_{i}\leq g_{i}^{T}g.$ (19)
$\mathrm{max}_{\lambda\geq 0}\mathrm{min}_{g}\;g^{T}Hg+\lambda(e_{i}-g_{i}^{T}g)=:L(g,\lambda).$ (20)
Then, the solution of the dual problem is obtained as $g^{*}(\lambda)=\mathrm{argmin}_{g}L(g,\lambda)=\frac{1}{2}\lambda H^{-1}g_{i}$. Using $g^{*}(\lambda)$, we can get the optimal Lagrange multiplier as $\lambda^{*}=\mathrm{argmax}_{\lambda}L(g^{*},\lambda)=2e_{i}/(g_{i}H^{-1}g_{i})=2\epsilon/\sqrt{g_{i}H^{-1}g_{i}}$. Then, the solution is $g^{*}=g^{*}(\lambda^{*})=\epsilon H^{-1}g_{i}/\sqrt{g_{i}^{T}H^{-1}g_{i}}$. As a result, the solutions of (5) and (6) are the same as $\epsilon H^{-1}g_{i}/\sqrt{g_{i}^{T}H^{-1}g_{i}}$.
A.2 Proof of Theorems
See 4.1

### v2

@@ 6 Experiments
This section evaluates the proposed method and baselines across various tasks with and without constraints. First, we explain how methods are evaluated on the tasks and then present the CMORL baselines. Subsequently, each task is described, and the results are analyzed. Finally, we conclude this section with ablation studies of the proposed method.
6.1 Evaluation Metrics and Baselines
Metrics. Once the universal policy is trained, an estimated CP front, denoted as $P\subset\mathbb{R}^{N}$, can be obtained by calculating the objective values $(J_{R_{1}}(\pi),...,J_{R_{N}}(\pi))$ for a range of preferences. Given that CMORL aims to approximate the ground truth CP front, it is essential to assess how close the estimated CP front is to the ground truth, and it can be realized by measuring the coverage and density of the estimated front. To this end, we use the hypervolume (HV) for coverage measurement and sparsity (SP) for density estimation, which are commonly used in many MORL algorithms [32, 7]. Given an estimated CP front $P\subset\mathbb{R}^{N}$ and a reference point $r\in\mathbb{R}^{N}$, the hypervolume (HV) is defined as follows [36]:
@@ 6.1 Evaluation Metrics and Baselines
where $Q(P,r):=\{z|\exists p\in P\;\text{s.t.}\;r\preceq z\preceq p\}$, and the metric represents the volume of the area surrounding the reference point and the estimated CP front. For a detailed explanation of this metric, including a visual description and how the reference point is set, please refer to Appendix C.2. Sparsity (SP), on the other hand, measures how uniformly the estimated CP fronts are distributed and is defined as the average of the squared distances between elements of the CP fronts [32]. However, the original definition exhibits a correlation with HV, where a larger HV corresponds to a larger SP. To eliminate this inherent correlation, we propose to use the following normalized version of SP:
$\overline{\mathrm{SP}}(P):=\frac{1}{|P|-1}\sum_{j=1}^{N}\sum_{i=1}^{|P|-1}\left(\frac{\tilde{P}_{j}[i]-\tilde{P}_{j}[i+1]}{\mathrm{max}_{k}\tilde{P}_{j}[k]-\mathrm{min}_{k}\tilde{P}_{j}[k]}\right)^{2},$ (14)
where $\tilde{P}_{j}:=\mathrm{Sort}(\{p[j]|\forall p\in P\})$, and $\tilde{P}_{j}[i]$ is the $i$th element in $\tilde{P}_{j}$. By normalizing the distance using the minimum and maximum values of the CP fronts, we can remove the correlation with HV while preserving the ability to measure sparsity.
Baselines. We aim to compare the proposed method with various existing CMORL algorithms; however, to the best of our knowledge, LP3 [17] is the only existing CMORL algorithm. To provide more comprehensive baselines, we include MORL algorithms that are based on policy gradient approaches, which can be adapted to handle constraints using the Lagrangian method. There are two state-of-the-art policy gradient-based MORL algorithms: PD-MORL [7], which is based on TD3 [14], and CAPQL [24], which is based on SAC [15]. We have extended them to CMORL by handling constraints using the Lagrangian method. For implementation details, please refer to Appendix C.3.1.
6.2 Legged Robot Locomotion
The legged robot locomotion tasks [20] are to control a quadrupedal or bipedal robot to follow randomly given commands while satisfying three constraints: keeping 1) body balance, 2) CoM height, and 3) pre-defined foot contact timing. Since the original tasks were designed to have a single objective, we have modified the tasks to have two objectives: 1) matching the current velocity with the command and 2) minimizing energy consumption. Please see Appendix C for more details.
The evaluation results are presented in Figure 4, and the estimated CP fronts are visualized in Appendix D. Across all tasks, CoMOGA achieves the highest HV and the lowest SP while satisfying all constraints. Notably, CoMOGA surpasses the baselines in the bipedal task. This suggests that preventing gradient conflicts is particularly effective for dynamically unstable robots, such as bipedal robots, because these robots require a careful balance of multiple objectives to attain desired behaviors. LP3 shows performance similar to the proposed method in the quadrupedal task but underperforms in the bipedal task. This may be attributed to its less effective handling of gradient conflicts.
Figure 4: Evaluation results of the legged robot locomotion tasks. The upper row shows results for the bipedal robot, while the lower row is for the quadrupedal robot. All algorithms are evaluated at every $10^{5}$ steps. The bold lines and shaded areas represent the mean and quarter-scaled standard deviation of results from five random seeds, respectively. The black dotted lines in constraint graphs indicate the thresholds.
6.3 Safety Gymnasium
We utilize single-agent and multi-agent goal tasks in the Safety Gymnasium environments [18]. These tasks require navigating robots to designated goals while avoiding obstacles. The single-agent goal tasks have two objectives: 1) reaching goals as many times as possible within a time limit and 2) maximizing energy efficiency, along with a single constraint to avoid collisions with obstacles. The multi-agent goal tasks have two objectives and two constraints, which are to maximize goal achievement and avoid collisions for two robots, respectively. In both tasks, point and car robots are used. For details on the tasks and hyperparameter settings, please refer to Appendix C.
The results are presented in Figure 5. CoMOGA shows the highest HV in all tasks and the lowest SP in two out of four tasks, while satisfying all constraints. Additionally, the low volatility in the evaluation graph indicates that the proposed method can stably handle both constraints and objectives. LP3 and PD-MORL violate constraints in at least one task, and CAPQL satisfies all constraints but shows weak performance in the single agent tasks, implying the difficulty of handling constraints and objectives simultaneously.
(a) Single-agent tasks. The left shows the results for the point robot, and the right is for the car robot.
(b) Multi-agent tasks. The upper row shows the results for the point robot, and the lower is for the car robot.
Figure 5: Evaluation results of the Safety Gymnasium tasks.
6.4 Multi-Objective Gymnasium
We conduct experiments in the Multi-Objective (MO) Gymnasium [4], which is a well-known MORL environment, to examine whether the proposed method also performs well on unconstrained MORL tasks. We use the MuJoCo tasks with continuous action spaces in the MO Gymnasium, and there are six tasks available: Hopper, Humanoid, Half-Cheetah, Walker2d, Ant, and Swimmer. Each task has three, two, two, two, three and two objectives, respectively. Details are provided in Appendix C.
The results are presented in Figure 6, and the proposed method achieves the highest HV in three tasks and the lowest SP in five out of six tasks. Especially, CoMOGA exhibits outstanding performance in the hopper and humanoid tasks, which suggests that the conflict-averse strategy is effective for dynamically unstable robots even in unconstrained MORL settings. However, all algorithms except LP3 show high HV results in the walker2d task, which is also dynamically unstable. It can be considered that they have reached the global optimum due to the well-designed reward functions.
Figure 6: Evaluation results of the MO Gymnasium tasks. Each column shows the results of the task corresponding to its title.
6.5 Ablation Study
Some might suggest that conflict aversion in CMORL can be achieved by simply integrating MTL algorithms with the Lagrangian method. To answer this question, we compare a method that combines a conflict-averse MTL algorithm, CAGrad [23], and the Lagrangian approach against CoMOGA in the legged locomotion tasks. The results are shown in Figure 7. CAGrad+Lagrangian violates the second constraint in the quadrupedal robot, and while it satisfies all constraints in the bipedal robot, HV and SP metrics are significantly lower than CoMOGA. These results demonstrate that the proposed method can handle both constraints and gradient conflicts effectively.
Figure 7: Results of the ablation study in the legged robot locomotion tasks.
7 Conclusions and Limitations
We have introduced a CMORL algorithm called CoMOGA to maximize multiple objectives while satisfying safety constraints. The proposed method is based on a novel transformation process that converts objectives into constraints, which enables the avoidance of gradient conflicts among multiple objectives and the handling of constraints without additional optimization variables. We show that the proposed method converges to a CP optimal policy in tabular settings and demonstrate that it achieves outstanding performance in HV and SP metrics with constraint satisfaction through various experiments. Since the proposed method updates the policy in a manner that prevents any objective functions from decreasing, the set of trained policies may cover a narrow portion of the CP front. Future research can address this limitation by mitigating the condition of gradient conflicts to better balance multiple objectives and expand coverage. Specifically, by satisfying the conditions of the proposed generalized policy update rule in Theorem 4.2, various CMORL algorithms can be developed with convergence guarantees, which can accelerate future work.
Appendix A Toy Example
Figure 8: Landscape of the objective functions of the toy example.
In this section, we describe the details of the toy example, which is initially defined in [23]. Objective functions of the toy example are formulated as follows:
$L_{1}(x_{1},x_{2})$ $:=c_{1}(x_{2})f_{1}(x_{1},x_{2})+c_{2}(x_{2})g_{1}(x_{1},x_{2}),$
$L_{2}(x_{1},x_{2})$ $:=c_{1}(x_{2})f_{2}(x_{1},x_{2})+c_{2}(x_{2})g_{2}(x_{1},x_{2}),\;\text{where}$
$c_{1}(x)$ $:=\max(\mathrm{tanh}(0.5x),0),$
$c_{2}(x)$ $:=\max(\mathrm{tanh}(-0.5x),0),$
$f_{1}(x_{1},x_{2})$ $:=\log(\max(|0.5(-x_{1}-7)-\mathrm{tanh}(-x_{2})|,0.000005))+6,$
$f_{2}(x_{1},x_{2})$ $:=\log(\max(|0.5(-x_{1}+3)+\mathrm{tanh}(-x_{2}+2)|,0.000005))+6,$
$g_{1}(x_{1},x_{2})$ $:=((-x_{1}+7)^{2}+0.1(-x_{2}-8)^{2})/10-20,$
$g_{2}(x_{1},x_{2})$ $:=((-x_{1}-7)^{2}+0.1(-x_{2}-8)^{2})/10-20,$
where $L_{1}$ and $L_{2}$ are the objective functions. The landscape of the objective functions are presented in Figure 8. In addition, we define a constraint function, which is defined as follows:
$C(x_{1},x_{2}):=x_{1}^{2}+0.3(x_{2}-10)^{2}-10.5^{2}.$
Finally, the toy example is formulated as follows:
$\min_{x_{1},x_{2}}L_{1}(x_{1},x_{2})\;\text{and}\;L_{2}(x_{1},x_{2})\quad\textbf{s.t.}\;C(x_{1},x_{2})\leq 0.$
We set the preference as $(0.5,0.5)$ for both algorithms, LS and CoMOGA, and use the following four initial points: $(-10,0)$, $(-10,7.5)$, $(0,7.5)$, and $(10,10)$. The optimization trajectories for each method are shown in Figure 2. For the initial point $(-10,7.5)$, the gradient of $L_{1}$ is much larger than $L_{2}$. Thus, addressing the toy example with LS can be significantly influenced by $L_{1}$ and eventually converges to a local optimal point of $L_{1}$. To resolve this issue, avoiding gradient conflicts can prevent policy gradients from being dominated by a specific objective, which helps them converge to an optimal point.
Appendix B Proofs
B.1 Proof of Transformation Process
In this section, we prove that the solutions of (2) and (3) are equivalent. Under the linear assumption, we can express the $i$th objective function as $g_{i}^{T}\Delta\theta+J_{R_{i}}(\theta_{\mathrm{old}})$. Then, (2) can be rewritten as follows:
$\max_{\Delta\theta}\;g_{i}^{T}\Delta\theta+J_{R_{i}}(\theta_{\mathrm{old}})\;\;\mathrm{s.t.}\;\Delta\theta^{T}H\Delta\theta\leq\epsilon^{2}.$
$\max_{\lambda\geq 0}\min_{\Delta\theta}\;-g_{i}^{T}\Delta\theta+\lambda(\Delta\theta^{T}H\Delta\theta-\epsilon^{2})=:L(\Delta\theta,\lambda).$
Then, $\Delta\theta^{*}(\lambda)=\mathrm{argmin}_{\Delta\theta}L(\Delta\theta,\lambda)=H^{-1}g_{i}/(2\lambda)$. By replacing $\Delta\theta$ with $\Delta\theta^{*}(\lambda)$ in the dual problem, we get $\lambda^{*}=\mathrm{argmax}_{\lambda}L(\Delta\theta^{*}(\lambda),\lambda)=\sqrt{g_{i}^{T}H^{-1}g_{i}}/(2\epsilon)$. Finally, we obtain the solution, $\Delta\theta^{*}=\Delta\theta^{*}(\lambda^{*})=\epsilon H^{-1}g_{i}/\sqrt{g_{i}^{T}H^{-1}g_{i}}$.
$\min_{\Delta\theta}\;\Delta\theta^{T}H\Delta\theta\;\;\mathrm{s.t.}\;e_{i}\leq g_{i}^{T}\Delta\theta.$ (15)
$\max_{\lambda\geq 0}\min_{\Delta\theta}\;\Delta\theta^{T}H\Delta\theta+\lambda(e_{i}-g_{i}^{T}\Delta\theta)=:L(\Delta\theta,\lambda).$ (16)
Then, the solution of the dual problem is obtained as $\Delta\theta^{*}(\lambda)=\mathrm{argmin}_{\Delta\theta}L(\Delta\theta,\lambda)=\lambda H^{-1}g_{i}/2$. Using $\Delta\theta^{*}(\lambda)$, we can get the optimal Lagrange multiplier as $\lambda^{*}=\mathrm{argmax}_{\lambda}L(\Delta\theta^{*}(\lambda),\lambda)=2e_{i}/(g_{i}H^{-1}g_{i})=2\epsilon/\sqrt{g_{i}H^{-1}g_{i}}$. Then, the solution is $\Delta\theta^{*}=\Delta\theta^{*}(\lambda^{*})=\epsilon H^{-1}g_{i}/\sqrt{g_{i}^{T}H^{-1}g_{i}}$. As a result, the solutions of (2) and (3) are the same as $\epsilon H^{-1}g_{i}/\sqrt{g_{i}^{T}H^{-1}g_{i}}$.
B.2 Proof of Theorem 4.2
We first define the following notions for a policy $\pi$:
$d_{\rho}^{\pi}(s)$ $:=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}\mathrm{Pr}(s_{t}=s)\Big{|}s_{0}\sim\rho,a_{t}\sim\pi(\cdot|s_{t}),s_{t+1}\sim P(\cdot|s_{t},a_{t})\;\forall t,$
$V^{\pi}(s)$ $:=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t},s_{t+1})\Big{|}s_{0}=s,a_{t}\sim\pi(\cdot|s_{t}),s_{t+1}\sim P(\cdot|s_{t},a_{t})\;\forall t\right],$
$Q^{\pi}(s,a)$ $:=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t},s_{t+1})\Big{|}s_{0}=s,a_{0}=a,s_{t+1}\sim P(\cdot|s_{t},a_{t}),a_{t+1}\sim\pi(\cdot|s_{t+1})\;\forall t\right],$
$A^{\pi}(s,a)$ $:=Q^{\pi}(s,a)-V^{\pi}(s).$
We also simplify several notations for brevity as follows:
$\pi_{\theta_{t}}\rightarrow\pi_{t},\;d_{\rho}^{\pi_{\theta_{t}}}\rightarrow d_{\rho}^{t},\;A_{R_{i}}^{\pi_{\theta_{t}}}(s,a)\rightarrow A_{R_{i}}^{t}(s,a),\;A_{C_{k}}^{\pi_{\theta_{t}}}(s,a)\rightarrow A_{C_{k}}^{t}(s,a).$
Since it is assumed that the state and action spaces are finite, a policy can be parameterized using a softmax parameterization as follows [3]:
$\pi_{\theta}(a|s):=\exp(\theta(s,a))\Big{/}\sum_{a^{\prime}\in A}\exp(\theta(s,a^{\prime})),$
where $\theta\in\mathbb{R}^{S\times A}$ is a trainable parameter, and $\theta(s,a)$ denotes the value corresponding to state $s$ and action $a$. Through Lemma 4 in [33], the following equations are satisfied when a policy is updated according to (9):
$\theta_{t+1}=\theta_{t}(s,a)+\frac{\alpha_{t}}{1-\gamma}\cdot\begin{cases}(\sum_{i=1}^{N}\nu_{t,i}^{a}A_{R_{i}}^{t}-\alpha_{t}\sum_{k=1}^{M}\lambda_{t,k}^{a}A_{C_{k}}^{t})&\text{if}\;J_{C_{k}}(\theta_{t})\leq d_{k}\;\forall k,\\ (\alpha_{t}\sum_{i=1}^{N}\nu_{t,i}^{b}A_{R_{i}}^{t}-\sum_{k=1}^{M}\lambda_{t,k}^{b}A_{C_{k}}^{t})&\text{otherwise},\end{cases}$
$\pi_{t+1}(a|s)=\pi_{t}(a|s)\frac{\exp(\theta_{t+1}(s,a)-\theta_{t}(s,a))}{Z_{t}(s)},$
$\text{where}\;Z_{t}(s):=\sum_{a\in A}\pi_{t}(a|s)\exp(\theta_{t+1}(s,a)-\theta_{t}(s,a)).$
Now, we introduce a lemma showing that the LS approach guarantees to converge a CP optimal policy, followed by the proof of Theorem 4.3.
Lemma B.1.
Given a preference $\omega$, let us define an LS optimal policy as follows:
@@ B.1 Proof of Transformation Process
Then, the LS optimal policy is also a CP optimal policy.